ABC Company sold 3000 units in October at a sales price of $45 per unit. The variable cost is $25 per unit. Calculate the total contribution margin also in percentage and per unit.
Answer #1

rate positively ..

Calculate the total contribution margin
Sales price = 45
Variable cost per unit = 25
Contribution margin 20
Total contribution margin = 60000
20*3000
Ans = 60000
contribution margin % = 60000/(45*3000) 44.44%
Ans = 44.44%
Per unit contribution margin = 45-25 = 20
ans = $   20.00 per unit